Louis and Carrie Anne are students who have set up a summer business in their neighborhood. They cut lawns and clean cars. Louis is particularly efficient at cutting the grasshe requires 1 and 1/4 hours to cut a typical lawn, while Carrie Anne needs 1 and 2/3 hours. In contrast, Carrie Anne can wash a car in 3/4 hour, while Louis requires 1 hour. (a) If they decide to specialize in the tasks, who should cut the grass and who should wash cars? (b) If they each work a ten hour day, how many lawns can they cut and how many cars can they wash if they each specialize in performing the task where they are most efficient? (c) Illustrate the PPF for each individual where lawns are on the horizontal axis and car washes on the vertical axis, if each individual has twelve hours in a day.


1
Expert's answer
2022-03-07T11:02:33-0500

a)

Louis has an advantage in cutting the grass while Carrie Anne should wash cars.

Reason.

Louis is faster by 25 minutes compared to Carrie Anne in cutting Lawns while Carrie Anne is faster by 15 minutes in washing cars compared to Louis.


b)

"=10 hrs\div \frac {5}{4}hrs"


"=8\space Lawns"


"=10hrs\div \frac {3}{4}hrs"


"=13.33"


"=13 \space Cars"


c)

Number of Lawns

"=12\div \frac {5}{4}hrs"


"=9\space Lawns"


Number of Cars

"=12\div \frac {3}{4}hrs"


"=16\space Cars"


Illustrate of the PPF